OpenStudy (anonymous):

Solve the inequality 7(x - 2) < -5. x < 9/7 x > 9/7 x < -9/7 x > -9/7

OpenStudy (anonymous):

Do you know what to do first

OpenStudy (anonymous):

no because the x's are already on one side

OpenStudy (anonymous):

7x-14<-5 then 7x<-5+14 7x<9 and x<9/7

OpenStudy (kendricklamar2014):

1. Simplify both sides:\[7x−14<−5\]

OpenStudy (kendricklamar2014):

2. Add 14 to both sides:\[7x−14+14<−5+14 = 7x<9\]

OpenStudy (kendricklamar2014):

3. Divide both sides by 7:\[\frac{ 7x }{ 7} < \frac{ 9 }{ 7 }\]

OpenStudy (kendricklamar2014):

Which give you an answer of:\[x < \frac{ 9 }{ 7 }\]

OpenStudy (anonymous):

thank you so much

OpenStudy (kendricklamar2014):

You're Welcome